OpenStudy (anonymous):

What is the slope of the line x = -5?

OpenStudy (amoodarya):

\[\infty\\\]

OpenStudy (amoodarya):

 x=-5 is parallel with y axis

OpenStudy (anonymous):

undefined?

OpenStudy (amoodarya):

 \[\lim_{\theta \rightarrow \pi/2}\tan \theta \rightarrow \]